OpenStudy (anonymous):

what is the mean, median, mode of data : a) 25,21,38,29,32,44,38,21,16 4:47 pm

OpenStudy (anonymous):

Whatcha need? I can help. You've got the median and mode already, yeah?

OpenStudy (anonymous):

yea but i need another help too

OpenStudy (anonymous):

25,21,38,29,32,44,38,21,16 Helps to put them in order first: 16, 21, 21, 25, 29, 32, 38, 38, 44 From here, the mean is the average. So take all of the numbers, add them up and divide them by the number of numbers you have in your list (in this case nine). Mode is the number that shows up the most. In this case you have two numbers that each show up twice. Median is the number in the middle of the sequence. Start from the outsides and cross em off one by one, one from the left, one from the right, until you reach the center number. This one is easy because there are an odd number of numbers, you will get one number for the median. Hope that helps.

OpenStudy (anonymous):

so the median would be 58?

OpenStudy (anonymous):

The median, or middle number is 29. The MEAN, which means average, without doing any math, I can tell you it is not 58. Add all the numbers together and then divide by the number of numbers in the sequence. There are nine numbers in this sequence. Add them and divide by nine.

OpenStudy (anonymous):

29

OpenStudy (anonymous):

29 and 1/3.

OpenStudy (anonymous):

(16 + 21 + 21 + 25 + 29 + 32 + 38 + 38 + 44)/9. 1 2 3 4 5 6 7 8 9

OpenStudy (anonymous):

sorry i gave the wrong number ..so for instance i rearranger the numbers in to median such as , 36,38,44,53,58,71,74,81,82,94 .. so the median would be 58 ??

OpenStudy (anonymous):

Also, I just read chat. There is such a thing as having 2 modes.

OpenStudy (anonymous):

the median is 58+71/2

OpenStudy (anonymous):

In that case, you have 10 numbers so there will be two numbers in the center. The median is those two numbers added and then divided by 2.

OpenStudy (anonymous):

yay i got it :) thankyou

OpenStudy (anonymous):

Np.

OpenStudy (anonymous):

Could you help me as well?

OpenStudy (anonymous):

I am almost out of time, but if it's in my knowledge, I will do my best to help.

OpenStudy (anonymous):

Here's the equation -> 4x^2+20x-24=0 and I've gotta factor and solve. I'm able to get the first step, and afterwards I dunno where to go.

OpenStudy (anonymous):

So let's take out 4 first. Notice how the coefficients and constant can all be divided by 4 so we can do: 4(x^2+5x-6)=0

OpenStudy (anonymous):

If you distrubute that, you'll notice that it is the same as 4x^2+20x-24=0

OpenStudy (anonymous):

I get that much.

OpenStudy (anonymous):

Awesome.

OpenStudy (anonymous):

We are going to leave that four alone for now. I'll explain it later. Can you factor (x^2+5x-6)?

OpenStudy (anonymous):

Remember, the formula looks like (x+?)(x+?).

OpenStudy (anonymous):

It'd be

OpenStudy (anonymous):

(x+6)(x-1)

OpenStudy (anonymous):

The two ? should ADD to 5, the MIDDLE term. The two ? should MULTIPLY to -6, the LAST TERM (aka the constant).

OpenStudy (anonymous):

Great, that is exactly right.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

So, if (x+6)(x-1) = (x^2+5x-6), can you agree that 4(x^2+5x-6)) = 4 (x+6)(x-1) ?

OpenStudy (anonymous):

So the four is allowed to stay out there ;) All remains balanced.

OpenStudy (anonymous):

Are we clear on this?

OpenStudy (anonymous):

the table shows the prizes awarded in jason's school funraiser : price value : 100,50,10 , number of prices 2,4,10 find the mean median mode

OpenStudy (anonymous):

Yeah. Do you have to go now?

OpenStudy (anonymous):

I have to change classes, but I have computers in there too ;) It might be a couple minutes before I can come back.

OpenStudy (anonymous):

The bell just rang, so I will be right back in maybe 10.

OpenStudy (anonymous):

Okay. Tell me later what age you are if you can.

OpenStudy (anonymous):

the table shows the prizes awarded in jason's school funraiser : price value : 100,50,10 , number of prices 2,4,10 find the mean median mode hellllpppp

OpenStudy (anonymous):

Okay, I can, give me a minute

OpenStudy (anonymous):

:) please

OpenStudy (anonymous):

Eh

OpenStudy (anonymous):

I'm confused by the number of prices part

OpenStudy (anonymous):

what part you go confused?

OpenStudy (anonymous):

Well, did you mean number of prizes, maybe?

OpenStudy (anonymous):

yes thats exactly what i meant

OpenStudy (anonymous):

Okay so, to solve for the three..beginning with median..

OpenStudy (anonymous):

yh?

OpenStudy (anonymous):

I think you needa arrange the numbers in order from lowest to highest.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

k for median i got 31.25 ,

OpenStudy (anonymous):

I believe you solved for the mean.

OpenStudy (anonymous):

Noooo i dont get mean and mode and thatss y i need your help

OpenStudy (anonymous):

Mkay, so for mean: you must add up all the values and divide by the number of values listed.

OpenStudy (anonymous):

The mode is the most common number, so it'd be 10.

OpenStudy (anonymous):

And then the median is what's in the middle: med=mid; you could remember it that way, I suppose.

OpenStudy (anonymous):

You have six values, which is an even number, so there isn't really a centered value, so in this case, you select the two values in the middle and take their average.

OpenStudy (anonymous):

i still dont get mean :S is it 8?

OpenStudy (anonymous):

10 + 50 + 10 + 2 + 4 + 10 = Your answer. So use a calculator to add those values up first.

OpenStudy (anonymous):

u missed 100 i guess

OpenStudy (anonymous):

My bad, typo, it would be 100, yeah.

OpenStudy (anonymous):

176

OpenStudy (anonymous):

/2?

OpenStudy (anonymous):

Good, now divide by how many values/numbers are listed.

OpenStudy (anonymous):

So it'd be six.

OpenStudy (anonymous):

29.3

OpenStudy (anonymous):

Correct. That's the mean. The average. Now you needa find the median.

OpenStudy (anonymous):

how am i gonaa??

OpenStudy (anonymous):

2, 4, 10, 10, 50, 100

OpenStudy (anonymous):

See how there isn't a number directly in the middle?

OpenStudy (anonymous):

So you'd either take the average of the two middle-ish numbers, and if they're the same, they're both considered the median.

OpenStudy (anonymous):

If it were 2, 4, 7, 9, 50, and 100, the median would be 9.

OpenStudy (anonymous):

Kidding, 8.

OpenStudy (anonymous):

yea i told u in the beh=ginig itss 8 lol but back of the book it shows 10

OpenStudy (anonymous):

Do you need to show your work? If you have answers in the back of the book, you just need to jot down the steps that give you those answers.

OpenStudy (anonymous):

i said back of the book its sayss 10 for the median.. but i got 8 .. so i dont kno whats the right answer

OpenStudy (anonymous):

Tell me how you got eight.

OpenStudy (anonymous):

i added the number of prices which is 2+4+6=16/2

OpenStudy (anonymous):

Median is middle, remember that. Med=mid.

OpenStudy (anonymous):

So you don't even need to do any calculating, you just need to determine which value/number is in the center.

OpenStudy (anonymous):

oh i got it :)

OpenStudy (anonymous):

And I hafta go soon, I have a math exam tomorrow morning that I need to study for, but I want you to solve for mean, median and mode of the example equation I'm gonna give you.

OpenStudy (anonymous):

Very nice guys, good deal. I got a little waylaid but everything looks to be in order ;)

OpenStudy (anonymous):

thankyou soo much :)good luck for ur exam

OpenStudy (anonymous):

Dude, IParke, can you make sure this fella knows what he's doing? I dunno if he's got it down yet, so if you've got time, give him random scenarios to work with ie finding mean, median and mode and all. Can you do that?

OpenStudy (anonymous):

If he needs help, I got him covered.

OpenStudy (anonymous):

MSW, would you like a practice set of numbers?

OpenStudy (anonymous):

hahahahhahhaa :P *this fellow *was referring to me ??? :P

OpenStudy (bahrom7893):

lol why the heck do ur questions have over 80 replies all the time??

OpenStudy (anonymous):

Please go away if you're not gonna assist. That's a rhetorical question, anyway.

OpenStudy (anonymous):

And yes MSW, I was referring to you.

OpenStudy (bahrom7893):

i do assist actually.. anyone needs help?

OpenStudy (anonymous):

lol !! well ... i felt bad

OpenStudy (anonymous):

Why's that?

OpenStudy (bahrom7893):

And Aubburrito please don't tell me what to do.

OpenStudy (anonymous):

coz im a girl :P lol hahah

OpenStudy (anonymous):

Fella goes for either gender, I believe.

OpenStudy (anonymous):

oh comonnnn :P

OpenStudy (anonymous):

That's why I chose it! Geeeeesh. Tough crowd.

OpenStudy (anonymous):

haha ! okay